OpenStudy (anonymous):

find an equation of the line having the given slope and containing the given point. m=4/5, (2,-9) The equation of the line in slope intercept form is y= ?

OpenStudy (anonymous):

@acxbox22

OpenStudy (acxbox22):

to find the y-intercept you have to plug in your values in the equation y=mx+b so you get 2=2(8)+b and then b=-14 so now you can find your equation which is y=2x-14